A particular medication dosage is 20.1 mg/kg of body weight. If 1.00 mL of the medication contains 50.0 mg, how many mL of the medication should a child weighing 59.0 lbs. receive?
1
Expert's answer
2019-09-16T04:03:55-0400

Convert lb in kg:

"59.0 lb\times\frac{1kg}{2.2046 lb}= 26.8 kg"


Find mg of medicine per 26.8 kg of body weight:

"\frac{20.1 mg}{1 kg}\times 26.8 kg = 538.68 mg"


Find volume of solution, containing 538.68 g of medicine:

"\frac{1ml}{50.0 mg}\times538.68 mg=10.8 ml"